# The Faiss Library

### v2

@@ Abstract
Vector databases typically manage large collections of embedding vectors. Currently, AI applications are growing rapidly, and so is the number of embeddings that need to be stored and indexed. The Faiss library is dedicated to vector similarity search, a core functionality of vector databases. Faiss is a toolkit of indexing methods and related primitives used to search, cluster, compress and transform vectors. This paper describes the trade-off space of vector search and the design principles of Faiss in terms of structure, approach to optimization and interfacing. We benchmark key features of the library and discuss a few selected applications to highlight its broad applicability.
@@ 1 Introduction
The emergence of deep learning has induced a shift in how complex data is stored and searched, noticeably by the development of embeddings. Embeddings are vector representations, typically produced by a neural network, that map (embed) the input media item into a vector space, where the locality encodes the semantics of the input. Embeddings are extracted from various forms of media: words [Mikolov et al., 2013, Bojanowski et al., 2017], text [Devlin et al., 2018, Izacard et al., 2021], images [Caron et al., 2021, Pizzi et al., 2022], users and items for recommendation [Paterek, 2007]. They can even encode object relations, for instance multi-modal text-image or text-audio relations [Duquenne et al., 2023, Radford et al., 2021].
Embeddings are employed as an intermediate representation for further processing, e.g. self-supervised image embeddings are input to shallow supervised image classifiers [Caron et al., 2018, Caron et al., 2021]. They are also leveraged as a pretext task for self-supervision [Chen et al., 2020]. In fact, embeddings are a compact intermediate representation that can be re-used for several purposes.
In this paper we consider embeddings used directly to compare media items. The embedding extractor is designed such that the distance between embeddings reflects the similarity between their corresponding media. As a result, conducting neighborhood search in this vector space offers a direct implementation of similarity search between media items.
@@
From a practical perspective, there are multiple advantages to maintain a clear separation of roles between the embedding extraction and the vector search algorithm. Both are bound by an “embedding contract” on the embedding distance:
  • The embedding extractor, which is typically a neural network in modern systems, is trained so that distances between embeddings are aligned with the task to perform.
@@
Faiss is a library for ANNS. The core library is a collection of source files written in standard C++ without dependencies. Faiss also provides a comprehensive Python wrapper for its C++ core. It is designed to be used both from simple scripts and as a building block of a DBMS. In contrast with other libraries that focus on a single indexing method, Faiss is a toolbox that contains a variety of indexing methods that commonly involve a chain of components (preprocessing, compression, non-exhaustive search, etc.). In this paper, we show that there exists a choice between a dozen index types, and the optimal one usually depends on the problem’s constraints.
To summarize what Faiss is not: Faiss does not extract features – it only indexes embeddings that have been extracted by a different mechanism; Faiss is not a service – it only provides functions that are run as part of the calling process on the local machine; Faiss is not a database – it does not provide concurrent write access, load balancing, sharding, transaction management or query optimization. The scope of the library is intentionally limited to focus on ANNS algorithmic implementation.
The basic structure of Faiss is an index that can have multiple implementations described in this paper. An index can store a number of database vectors that are progressively added to it. At search time, a query vector is submitted to the index. The index returns the database vector that is closest to the query vector in terms of Euclidean distance. There are many variants of this basic functionality: instead of just the nearest neighbor, $k$ nearest neighbors can be returned; instead of a fixed number of neighbors, only the vectors within a certain range can be returned; batches of vectors can be searched in parallel; other metrics besides the Euclidean distance are supported; the search can use either CPUs or GPUs.
Since it was open-sourced in 2017, Faiss became one of the most popular vector search libraries, totaling 30k GitHub stars. The Faiss packages have been downloaded 3M times. So far, 3800 works cite [Johnson et al., 2019], the reference paper for Faiss. Major vector database companies, such as Zilliz and Pinecone, either rely on Faiss as their core engine or have reimplemented Faiss algorithms.
This paper exposes the design principles of Faiss. A similarity search library has to trade off between different constraints (Section 3) using two main tools: vector compression (Section 4) and non-exhaustive search (Section 5). Faiss is engineered to be flexible and usable from other tools (Section 7). We also review a few applications of Faiss for trillion-scale indexing, text retrieval, data mining, and content moderation (Section 8). Throughout the paper, we refer to functions or classes in the Faiss codebase [^1] https://github.com/facebookresearch/faiss as well as the documentation [^2] https://faiss.ai/ using this style .
@@ 2 Related work
In the last decade, there has been a steady stream of papers about indexing methods that were published together with their reference implementations. In Faiss, we consider algorithms that cover a wide spectrum of use-cases.
One of the most popular approaches in the industry is to employ Locality Sensitive Hashing (LSH) as a way to compress embeddings into compact codes. In particular, the Cosine sketch [Charikar, 2002] produces binary vectors such that, in the corresponding Hamming space, the Hamming distance is an estimator of the cosine similarity between the original embeddings. The compactness of these sketches enables storing and therefore searching very large databases of media content [Lv et al., 2004], without the requirement to store the original embeddings. We refer the reader the early survey by [Wang et al., 2015] for research on binary codes.
Since the work by [Jégou et al., 2010], ANN based on quantization has emerged as a powerful alternative to binary codes [Wang et al., 2017]. We refer the reader to the survey by [Matsui et al., 2018] that discusses numerous research works related to quantization-based compact codes.
LSH often refers to indexing with multiple partitions, such as E2LSH [Datar et al., 2004]. We do no consider such scenario since the performance is not as good as using learnt partitions on real data [Paulevé et al., 2010]. Early data-aware methods that proved successful on large datasets include multiple partitions based on kd-tree or hierarchical k-means in [Muja and Lowe, 2014]. They are often combined with compressed-domain representation and are especially appropriate for very large-scale settings [Jegou et al., 2008, Jégou et al., 2010].
After the introduction of the NN-descent algorithm [Dong et al., 2011], ANN algorithms based on graphs have emerged as a viable alternative to methods based on space partitioning. In particular HNSW, which is the most popular current indexing method [Malkov and Yashunin, 2018] for medium-sized dataset is implemented in HNSWlib.
@@
Most of the research works on vector search have been open-sourced, and some of these evolved in relatively comprehensive software packages for vector search. For instance, FLANN includes several index types and a distributed implementation described extensively in [Muja and Lowe, 2014]. The first implementation of product quantization relied on the Yael library [Douze and Jégou, 2014], that already had a few of the Faiss principles: optimized primitives for clustering methods (GMM and k-means), scripting language interface (Matlab and Python) and benchmarking operators. The NMSlib package, intended for text retrieval was the first package to include HNSW [Boytsov et al., 2016] and also offers several index types. The HNSWlib library later became the reference implementation of HNSW [Malkov and Yashunin, 2018]. Google’s SCANN library is mainly a thoroughly optimized implementation of IVFPQ [Jégou et al., 2010] on SIMD and includes several index variants for various database scales. SCANN was open-sourced together with [Guo et al., 2020], which does not describe the engineering optimizations that make the library so fast. Diskann [Subramanya et al., 2019] is Microsoft’s foundational graph-based vector search library, initially built to exploit hybrid RAM/flash memory, but that also offers a RAM-only version called Vamana. It was later extended to perform efficient updates [Singh et al., 2021], out-of-distribution search [Jaiswal et al., 2022] and filtered search [Gollapudi et al., 2023].
Faiss was open-sourced simultaneously with the release of [Johnson et al., 2019], this publication describes the GPU implementation of several index types. The present paper complements this previous work by describing the library as a whole.
In parallel, numerous software libraries from the database world were extended or developed to do vector search. Milvus [Wang et al., 2021] uses its Knowhere library, which relies on Faiss as one of its core engines. Pinecone [Bruch et al., 2023] initially relied on Faiss. The engine has since been rewritten in Rust. Weaviate [van Luijt and Verhagen, 2020] is a composite retrieval engine that includes vector search.
@@
The leading benchmark for million-scale datasets is ANN-benchmarks [Aumüller et al., 2020] that compares about 50 implementations of ANNS. This benchmark was upgraded with the big-ANN [Simhadri et al., 2022a] challenge, which includes 6 datasets containing 1 billion vectors each. Faiss was used as a baseline for the challenge and multiple submissions were derived from it. The 2023 edition of the challenge is at a smaller scale (10M vectors) but the tasks are more elaborate. For instance, there is a filtered track for which Faiss was a baseline method [big,].
@@
Early datasets are based on keypoint features like SIFT [Lowe, 2004] used in image matching. We use BIGANN [Jégou et al., 2011b], a dataset of 128-dimensional SIFT features. Later, when global image descriptors produced by neural nets became popular, the Deep1B dataset was released [Babenko and Lempitsky, 2016], with 96-dimensional image features extracted with Google LeNet [Szegedy et al., 2015]. For this paper we introduce a dataset of 768-dimensional Contriever text embeddings [Izacard et al., 2021] that are compared using inner product similarity. The embeddings are computed with English Wikipedia passages. The higher dimension of these embeddings is typical for contemporary applications.
Each dataset has 10k query vectors, 20M to 350M training vectors. We indicate the size of the database explicitly, for example “Deep1M” means the database contains the 1M first vectors of deep1B. The training, database and query vectors are sampled randomly from the same distribution: we do not address out-of-distribution data [Jaiswal et al., 2022, Baranchuk et al., 2023] in this work.
@@ 3 Performance axes of a vector search library
$n=\underset{i=1..N}{\mathrm{argmin}}\|q-x_{i}\|$ (1)
@@
$(n_{1},...,n_{k},*,...,*)=\underset{i=1..N}{\mathrm{argsort}}\|q-x_{i}\|$ (2)
@@
These measures can be made equivalent by preprocessing transformations on the query and/or the database vectors. Table 2 summarizes the preprocessing transformations that are applicable for various bridges. Some were already identified [Bachrach et al., 2014, Hong et al., 2019], others are new.
Note that vectors transformed in this way have a very anisotropic distribution [Morozov and Babenko, 2018] and can be “harder” to index. In particular, for product or scalar quantization, the additional dimension incurred for many transformations is not homogeneous with other dimensions. See Section 4.2 for mitigations.
@@ 3.1 Brute force search
Implementing brute force search efficiently is not trivial [Chern et al., 2022, Johnson et al., 2019]. It requires (1) an efficient way of computing the distances and (2) for k-nearest neighbor search, an efficient way of keeping track of the $k$ smallest distances.
Computing distances in Faiss is performed either by direct distance computations, or, when query vectors are provided in large enough batches, using a matrix multiplication decomposition [Johnson et al., 2019, equation 2]. The corresponding Faiss functions are exposed in knn and knn_gpu for CPU and GPU respectively.
Collecting the top-$k$ smallest distances is usually done via a binary heap on CPU [Douze and Jégou, 2014, section 2.1] or a sorting network on GPU [Johnson et al., 2019, Ootomo et al., 2023]. For larger values of $k$, it is more efficient to use a reservoir: an unordered result buffer of size $k^{\prime}>k$ that is resized to $k$ when it overflows.
The Faiss IndexFlat implements brute force search. However, for large datasets this approach becomes too slow. In low dimensions, there are branch-and-bound methods that yield exact search results. However, in large dimensions they provide no speedup over brute force search [Weber et al., 1998].
Besides, for some applications, small variations in distances are not significant to distinguish application-level positive and negative items [Szilvasy et al., 2024]. In these cases, approximate nearest neighbor search (ANNS) becomes interesting.
@@ 3.2 Metrics for Approximate Nearest Neighbor Search
With ANNS, the user accepts imperfect results, which opens the door to a new solution design space. The database may be preprocessed into an indexing structure, rather than just being stored as a plain matrix.
@@
In ANNS the accuracy is measured as a discrepancy with the exact search results from (2) and (3). Note that this is an intermediate goal: the end-to-end accuracy depends on (1) how well the distance metric correlates with the item matching objective and (2) the quality of ANNS, which is what we measure here.
The accuracy for k-nearest neighbor search is generally evaluated as the “$n$-recall@$k$”, which is the fraction of the $n$ ground-truth nearest neighbors that are in the $k$ first search results (with $n\leq k$). Most often $k$$=$$1$ or $n$$=$$k$ (in which case the measure is a.k.a. “intersection measure”). When $n$ = $k$ = $1$, the recall measure and intersection are the same, and the recall is called “accuracy”. Note, in some publications [Jégou et al., 2010], recall@$k$ means 1-recall@$k$, while in others [Simhadri et al., 2022b] it corresponds to $k$-recall@$k$.
For range search, the exact search result is obtained by applying (3), using threshold $\varepsilon$. To yield result list $\widehat{R}$, the approximate search uses a (possibly different) threshold $\varepsilon^{\prime}$. Thus, standard retrieval metrics can be computed: precision $P=|R\cap\widehat{R}|/|\widehat{R}|$ and recall $R=|R\cap\widehat{R}|/|R|$. By sweeping $\varepsilon^{\prime}$ from small to large, the result list $\widehat{R}$ increases, producing a precision-recall curve. The area under the PR-curve is the mean average precision score of range search [Simhadri et al., 2022a]. Having $\varepsilon^{\prime}\neq\varepsilon$ makes sense when approximate vector representations distort the distance metric (more on this in Section 4).
For vector encoder-decoder pairs, the standard metric is the mean squared error (MSE) between the original vector and the reconstructed vector [Jégou et al., 2010, Babenko and Lempitsky, 2014, Huijben et al., 2024]. For an encoder $C$ and a decoder $D$, the MSE is:
$\mathrm{MSE}=\mathbb{E}_{x}\big{[}\|D(C(x))-x\|_{2}^{2}\big{]}$ (4)
@@
The other axes of the tradeoff are related to computing resources. During search, the search time and memory usage are the main constraints. If compression is used the memory usage can be smaller than that of the original vectors.
The index may need to store training data, which incurs a constant memory overhead before any vector is added to the index. The index can also add per-vector memory overhead to the memory used to store each vector. This is the case for graph indexes, that need to store graph edges for each vector. The memory usage is more complex for settings with hybrid storage (small fast memory + large slow memory), for example RAM + flash or GPU memory + RAM.
The index building time is also a resource constraint. It may be decomposed into a training time, independent of the number of vectors added to the index and the addition time per vector.
In distributed settings or flash-backed storage, the relevant metric is the number of I/O operations (IOPS) becasue every read operation fetches a whole page (of e.g. 4 kiB). It is better to lay out data to minimize the IOPs [Subramanya et al., 2019] than to perform small random accesses. Another possibly limiting factor is the amount of extra memory needed to pre-compute lookup tables used to speed up search.

### v3

@@ Abstract
Vector databases typically manage large collections of embedding vectors. Currently, AI applications are growing rapidly, consequently, the number of embeddings that need to be stored and indexed is increasing. The Faiss library is dedicated to vector similarity search, a core functionality of vector databases. Faiss is a toolkit of indexing methods and related primitives used to search, cluster, compress and transform vectors. This paper describes the trade-offs in vector search and the design principles of Faiss in terms of structure, approach to optimization and interfacing. We benchmark key features of the library and discuss a few selected applications to highlight its broad applicability.
@@ 1 Introduction
The emergence of deep learning has induced a shift in how complex data is stored and searched, noticeably by the development of embeddings. Embeddings are vector representations, typically produced by a neural network, that map (embed) the input media item into a vector space, where the locality encodes the semantics of the input. Embeddings are extracted from various forms of media: words [59, 10], text [24, 40], images [15, 69], users and items for recommendation [65]. They can even encode object relations, for instance multi-modal text-image or text-audio relations [31, 70].
Embeddings are employed as an intermediate representation for further processing, e.g. self-supervised image embeddings are input to shallow supervised image classifiers [14, 15]. They are also leveraged as a pretext task for self-supervision [18]. In fact, embeddings are a compact intermediate representation that can be re-used for several purposes.
In this paper, we consider embeddings used directly to compare media items. The embedding extractor is designed so that the distance between embeddings reflects the similarity between their corresponding media. As a result, conducting neighborhood search in this vector space offers a direct implementation of similarity search between media items.
@@
From a practical perspective, the embedding extraction and the vector search algorithm are bound by an “embedding contract” on the embedding distance:
  • The embedding extractor, typically a neural network in modern systems, is trained so that distances between embeddings are aligned with the task to perform.
@@
Faiss is a library for ANNS. The core library is a collection of C++ source files without external dependencies. Faiss also provides a comprehensive Python wrapper for its C++ core. It is designed to be used both from simple scripts and as a building block of a DBMS. In contrast with other libraries that focus on a single indexing method, Faiss is a toolbox that contains a variety of indexing methods that commonly involve a chain of components (preprocessing, compression, non-exhaustive search, etc.). In this paper, we show that there exists a choice between a dozen index types, and the optimal one usually depends on the problem’s constraints.
To summarize what Faiss is not: Faiss does not extract features – it only indexes embeddings that have been extracted by a different mechanism; Faiss is not a service – it only provides functions that are run as part of the calling process on the local machine; Faiss is not a database – it does not provide concurrent write access, load balancing, sharding, transaction management or query optimization. The scope of the library is intentionally limited to ANNS algorithmic implementation.
The basic structure of Faiss is an index that can have multiple implementations described in this paper. An index can store a number of database vectors that are progressively added to it. At search time, a query vector is submitted to the index. The index returns the database vector that is closest to the query vector w.r.t. the Euclidean distance. There are many variants of this functionality: instead of just the nearest neighbor, $k$ nearest neighbors are returned; instead of a fixed number of neighbors, only the vectors within a certain range are returned; batches of vectors can be searched in parallel; other metrics besides the Euclidean distance are supported; the search can use either CPUs or GPUs.
Since its open-source release in 2017, Faiss has emerged as one of the most popular vector search libraries, boasting 30k GitHub stars and more than 4000 citations of its GPU implementation paper [47]. The Faiss packages have been downloaded 3M times. Major vector database companies, such as Zilliz and Pinecone, either rely on Faiss as their core engine or have reimplemented Faiss algorithms.
This paper exposes the design principles of Faiss. A similarity search library has to trade off between different constraints (Section 3) using two main tools: vector compression (Section 4) and non-exhaustive search (Section 5). We also review a few applications of Faiss for trillion-scale indexing, text retrieval, data mining, and content moderation (Section 7). The appendix discusses how Faiss is structured and engineered to be flexible and usable from other tools. Throughout the paper, we refer to functions or classes in the Faiss codebase [^1] https://github.com/facebookresearch/faiss as well as the documentation [^2] https://faiss.ai/ using this specific style.
@@ 2 Related work
Over the past decade, a steady stream of papers about indexing methods has been published. Faiss encompasses a broad range of algorithms, catering to a diverse spectrum of use cases.
One of the most popular approaches in the industry is to employ Locality Sensitive Hashing (LSH) as a way to compress embeddings into compact codes. In particular, the cosine sketch [16] produces binary vectors such that the Hamming distance is an estimator of the cosine similarity between the original embeddings. The compactness of these sketches enables storing and searching very large databases of media content [54], without the requirement to store the original embeddings. We refer the reader to the early survey by [90] for research on binary codes.
Since the work by [44], quantization-based ANN has emerged as a powerful alternative to binary codes [89]. We refer the reader to the survey by [58] that discusses numerous research works related to quantization-based compact codes.
LSH often refers to indexing with multiple partitions, such as E2LSH [22]. However, we do not consider this scenario as the performance is generally inferior to that of learnt partitions [66]. Early data-aware methods that proved successful on large datasets include multiple partitions based on kd-tree or hierarchical k-means [62]. They are often combined with compressed-domain representation and are especially appropriate for very large-scale settings [43, 44].
After the introduction of NN-descent [26], graph-based ANN algorithms have emerged as a viable alternative to space partitioning methods. Notably, HNSW, currently the most popular indexing method [55] for medium-sized dataset, has been implemented in HNSWlib.
@@
Most of the research works on vector search have been open-sourced, and some of these evolved in relatively comprehensive software packages for vector search. For instance, FLANN includes several index types and a distributed implementation described extensively in [62]. The first implementation of product quantization relied on the Yael library [27], that already had a few of the Faiss principles: optimized primitives for clustering methods (GMM and k-means), scripting language interface (Matlab and Python) and benchmarking operators. NMSlib, a package originally designed for text retrieval, was the first to include HNSW [11]. It also provides several index types. The HNSWlib library later became the reference implementation of HNSW [55]. Google’s SCANN library is a thoroughly optimized implementation of IVFPQ [44] on SIMD and includes several index variants for various database scales. SCANN was open-sourced together with [36], which notably omits the engineering optimizations that underpin the library’s remarkable speed. Diskann [79] is Microsoft’s foundational graph-based vector search library, originally built to leverage hybrid RAM/flash memory. It also offers a RAM-only version. It was later extended to perform efficient updates [78], out-of-distribution search [42] and filtered search [34].
Faiss was open-sourced concurrently with the publication [47], which details the GPU implementation of several index types. The present paper complements this previous work by describing the library as a whole.
Concurrently, various software libraries from the database world were extended or developed to do vector search. Milvus [88] uses its Knowhere library, which relies on Faiss as one of its core engines. Pinecone [12] initially relied on Faiss, although the engine was later rewritten in Rust. Weaviate [86] is a composite retrieval engine that includes vector search among other methods.
Theoretical guarantees.
Several approximate algorithms were inspired by the Johnson-Lindenstrauss lemma, which stipulates that one can embed a set of high-dimensional points into a lower-dimensional space while almost preserving the distances. For instance, under certain assumptions about the data distribution, LSH algorithms based on random partitioning (e.g. projection) provide statistical guarantees for the range search problem. As this objective is only a proxy to the problem of nearest neighbor search, there are no formal guarantees w.r.t. any metrics such as nearest-neighbor recall. In practice, algorithms based on random partitioning are significantly outperformed by data-aware partitioning [66, 2]. Similarly, graph-based algorithms are difficult to compare from a theoretical perspective if the data distribution is unknown.
Techniques based on vector compression, either with binary sketches or quantization, often correspond to estimators of the target distance. For some of these methods, the variance of these estimators can be computed in closed form [16, 44], characterizing the distance estimation error. Again, the quality of the estimator is only a proxy for the actual k-nearest-neighbor problem.
Considering these factors and the importance of execution speed, the quality of ANN search is typically evaluated by experimental comparisons on publicly available benchmarks.
@@
The leading benchmark for million-scale datasets is ANN-benchmarks [3]. It compares about 50 implementations of ANNS. The big-ANN [76] challenge introduced large-scale settings, with 6 datasets containing 1 billion vectors each. Faiss was used as a baseline for the challenge and multiple submissions were derived from it. The 2023 edition of the challenge is at a smaller scale (10M vectors) but introduces more complex tasks, including a filtered track where Faiss served as a baseline method [75].
@@
Early datasets are based on keypoint features like SIFT [53] used in image matching. We use BIGANN [46], a dataset of 128-dimensional SIFT features. Later, when global image descriptors produced by neural nets became popular, the Deep1B dataset was released [6], with 96-dimensional image features extracted with Google LeNet [82]. For this paper, we introduce a dataset of 768-dimensional Contriever text embeddings [40] that are compared using inner product similarity. The embeddings are computed with English Wikipedia passages. The higher dimension of these embeddings is representative of contemporary applications.
Each dataset has 10k query vectors and 20M to 350M training vectors. We indicate the size of the database explicitly, for example “Deep1M” means the database contains the 1M first vectors of Deep1B. The training, database and query vectors are sampled randomly from the same distribution: we do not address out-of-distribution data [42, 8] in this work.
@@ 3 Performance axes of a vector search library
$n=\underset{i=1..N}{\mathrm{argmin}}\|q-x_{i}\|.$ (1)
@@
$(n_{1},...,n_{k},*,...,*)=\underset{i=1..N}{\mathrm{argsort}}\|q-x_{i}\|,$ (2)
@@
The Euclidean distance is the default for comparing vectors [53]. When vectors are obtained by metric learning, the cosine similarity is often used to avoid collapsing or exploding embeddings [23]. Maximum inner product search (MIPS) is most often used for recommendation systems that require to compare user and item embeddings [65].
These measures can be made equivalent by preprocessing transformations on the query and/or the database vectors. Table 2 summarizes the preprocessing transformations mapping different measures. To our knowledge, some of these were already identified [7, 37], while others are new.
Note that vectors transformed in this manner have a very anisotropic distribution [60]: the additional dimension incurred for many transformations is not homogeneous w.r.t. other dimensions. This can make indexing more difficult, in particular using product or scalar quantization methods. See Section 4.2 for mitigations.
@@ 3.1 Brute force search
Developing an efficient implementation of brute force search is not trivial [21, 47]. It requires (1) an efficient way of computing the distances and (2) an efficient way of keeping track of the $k$ smallest distances.
Computing distances in Faiss is performed either by direct distance computations, or, when query vectors are provided in large enough batches, using a matrix multiplication decomposition [47, Equation 2]. The corresponding Faiss functions are exposed in knn and knn_gpu for CPU and GPU, respectively.
Collecting the top-$k$ smallest distances is usually done via a binary heap on CPU [27, section 2.1] or a sorting network on GPU [47, 63]. For larger values of $k$, it is more efficient to use a reservoir: an unordered result buffer of size $k^{\prime}>k$ that is resized to $k$ when it overflows.
Faiss’s IndexFlat implements brute force search. However, for large datasets this approach becomes too slow. In low dimensions, there are branch-and-bound methods that yield exact search results. However, in large dimensions they provide no speedup over brute force search [91].
Besides, for some applications, small variations in distances are not significant enough to distinguish application-level positive and negative items [83]. In these cases, approximate nearest neighbor search (ANNS) becomes interesting.
@@ 3.2 Metrics for Approximate Nearest Neighbor Search
With ANNS, the user accepts imperfect results, which opens the door to a new solution design space. With exact search, the database is represented as a plain matrix. For ANNS, the database may be preprocessed into an indexing structure, more simply referred to as an index in the following.
@@
In ANNS the accuracy [^3] Metrics involved in a tradeoff are indicated in a specific font. is measured as a discrepancy with the exact search results from (2) and (3). Note that this is an intermediate goal: the end-to-end accuracy depends on (1) how well the distance metric correlates with the item matching objective and (2) the quality of ANNS, which is what we measure here.
Accuracy of k-nearest neighbor search is generally evaluated with “$n$-recall@$k$”, which is the fraction of the $n$ ground-truth nearest neighbors that are in the $k$ first search results (with $n\leq k$). Most often $k$$=$$1$ or $n$$=$$k$ (in which case the measure is a.k.a. “intersection measure”). When $n$ = $k$ = $1$, the recall measure and intersection are the same, and the recall is called “accuracy”. In some publications [44], recall@$k$ means 1-recall@$k$, while in others [77] it corresponds to $k$-recall@$k$.
For range search, the exact search result is obtained by applying (3), using threshold $\varepsilon$. To yield result list $\widehat{R}$, the approximate search uses a (possibly different) threshold $\varepsilon^{\prime}$. Thus, standard retrieval metrics can be computed: precision $P=|R\cap\widehat{R}|/|\widehat{R}|$ and recall $R=|R\cap\widehat{R}|/|R|$. By sweeping $\varepsilon^{\prime}$ from small to large, the result list $\widehat{R}$ increases, producing a precision-recall curve. The area under the PR-curve is the mean average precision score of range search [76]. Setting $\varepsilon^{\prime}\neq\varepsilon$ is relevant when approximate vector representations distort the distance metric (more on this in Section 4).
For vector encoder-decoder pairs, the standard metric is the mean squared error (MSE) between the original vector and the reconstructed vector [44, 4, 39]. For an encoder $C$ and a decoder $D$, the MSE is:
$\mathrm{MSE}=\mathbb{E}_{x}\big{[}\|D(C(x))-x\|_{2}^{2}\big{]}.$ (4)
@@
The other axes of the trade-off are related to computing resources. During search, the search time and memory usage are the main constraints. If compression is used, then, the memory usage can be smaller than what is required to store the original vectors.
The index may need to store training data, which incurs a constant memory overhead before any vector is added to the index. The index can also add per-vector memory overhead to the memory used to store each vector. This is the case for graph indexes, that need to store graph edges for each vector. The memory usage is more complex for settings with hybrid storage such as RAM + flash or GPU memory + RAM.
The index building time is also a resource constraint. It may be decomposed into a training time, which is independent of the number of vectors added to the index, and the addition time per vector.
In distributed settings or flash-backed storage, the relevant metric is the number of I/O operations (IOPS), as each read operation fetches a whole page. Data layouts that minimize IOPs [79] are more efficient than small random accesses. Another possibly limiting factor is the amount of extra memory needed to pre-compute lookup tables used to speed up search.